# Benchmarking Computer Science Survey Generation

### v1

@@ 7.1 Retrieval-Augmented Generation
Large Language Models (LLMs) are inherently limited by their static, pre-trained parametric knowledge. To address these limitations, Retrieval-Augmented Generation (RAG) has emerged as a key paradigm gao2023retrieval ; lewis2020retrieval ; dong2025decoupling ; tu2025rbft ; su2025dynamic . By grounding the model in external knowledge, RAG directly addresses several fundamental limitations of LLMs, offering a robust mechanism to mitigate hallucinations ji2023survey ; su2024unsupervised ; su2024mitigating ; wang2025joint , facilitate knowledge updating fang2024alphaedit ; wang2024lekube ; wang2024knowledge ; wang2025decoupling , and enable effective domain adaptation yang2024tcm ; su2025judge ; su2024stard ; su2024legalaid .
The conventional approach to traditional RAG is built upon the "Retrieval-then-Read" paradigm borgeaud2022improving ; guu2020retrieval ; lewis2020retrieval . Within this framework, a user’s query triggers a search module for relevant documents within a large-scale external corpus. This retrieval step is carried out by either an external retriever zhai2008statistical ; su2023caseformer ; robertson2009probabilistic ; su2023wikiformer ; ma2023caseencoder ; fang2024scaling or a more sophisticated retrieval system su2023thuir2 ; salemi2024towards . Building upon this foundation, recent work has proposed more advanced RAG architectures to improve efficiency and effectiveness. For instance, Dynamic RAG jiang2022retrieval ; su2024dragin ; yao2024seakr moves beyond a single retrieval step by adaptively triggering the retriever during generation, specifically when the LLM is uncertain during the generation process. From another angle, GraphRAG edge2024local enhances the knowledge source by querying pre-constructed knowledge graphs instead of unstructured text, allowing it to retrieve interconnected facts and relationships. Furthermore, the Parametric RAG paradigm  su2025parametric ; tan2025dynamic ; fleshman2025lora alters the knowledge injection step by directly injecting retrieved knowledge into the LLM’s parameters.
@@ 7.2 Long-Form Text Generation and Evaluation
Long-form text generation is substantially more challenging than short-text generation due to its inherent requirements for sustained coherence and rich contextual understanding. Early approaches mainly used generative adversarial networks and reinforcement learning to conduct long-sequence generation (Guo2018, ). More recently, large language models have emerged as a strong tool for this task, offering advanced capabilities to handle long-text generation. For example, structured planning techniques and specialized inference mechanisms are proposed (Sloan2025, ; Jin2024, ) to generate consistent and high-quality clinical reports. Similarly, hierarchical planning frameworks have demonstrated that content control and multi-constraint instruction following can significantly enhance logical flow and topic coverage (hu2022, ; pham2024, ). Beyond medical or other task-specific applications, context-driven retrieval strategies, such as tree-structured retrieval, can support open-domain long-text generation by guiding the model through extensive knowledge sources (roy2024, ). The effective evaluation framework is vital for measuring the quality, factualness, and user-centric utility of the long-form text generation task. Traditional metrics, designed for shorter texts, often fail to capture the intricacies of longer outputs. Recent work has introduced task-focused benchmarks that emphasize user-oriented objectives, such as personalized writing or domain-specific content generation (kumar2024, ; salemi2025, ). In parallel, factuality assessment has attracted growing interest, with methods proposed to evaluate both verifiable and unverifiable claims. Metrics such as VERISCORE and FACTSCORE break down generated text into atomic facts, checking each for consistency against reliable sources (song2024, ; min2023, ). Beyond factual correctness, coherence and structural quality have been studied extensively. Benchmarks like LongGenBench and HelloBench underscore the importance of evaluating a model’s ability to maintain logical organization and clarity over extended passages (wu2025, ; que2024, ).
@@ 7.3 Survey Generation
In the domain of scientific writing, survey generation involves distilling extensive textual resources into a coherent and structured overview. Recent advances in AI-assisted systems have provided prompting-based approaches to expedite the drafting process while preserving content accuracy (gero2022, ; Kacena2024, ). One of the most commonly used approaches is retrieval-augmented generation, which combines large-scale knowledge retrieved from documents with language generation empowered by LLMs to yield factually comprehensive overviews (lewis2021, ). Retrieval-augmented generation is often initiated with dense retrieval methods based on dual-encoder architectures to identify highly relevant documents (karpukhin2020, ). Once these documents are retrieved, summarization techniques—spanning top-down, bottom-up, and graph-based ranking methods—play a pivotal role in producing concise yet faithful summaries (nayeem2024, ; pang2022, ; bleiweiss2023, ). Building on these retrieval and summarization-based methodologies, automated literature survey generation has garnered increasing attention (nlpcc2024, ; lai2024, ). However, existing techniques depend on limited ground truths and employ coarse evaluation metrics, resulting in oversimplified assessments of survey quality wang2024autosurvey . To address these challenges, we present a refined ground truth and a multi-dimensional evaluation framework that emphasizes both accuracy and structural coherence. By evaluating quality through multiple dimensions, our proposed framework advances the capabilities of automated survey generation, offering a more comprehensive and rigorous approach to summarizing scientific literature.
@@ 8 Conclusion
In this paper, we introduce SurGE, a benchmark for end-to-end scientific survey generation in computer science. We construct a high-quality dataset and develop a multi-dimensional evaluation framework that assesses the coverage, coherence, factual accuracy, and writing quality of automatically generated surveys. Extensive experiments with various baselines highlight the challenges in producing high-quality surveys. SurGE is expected to serve as a valuable resource for future research in information retrieval, fostering advancements in both retrieval methodologies and generative AI techniques. Several limitations guide future work, including: (1) The benchmark is limited in the domain of computer science but could be extended to broader domains. (2)  Integration with more advanced retrieval and generation methods could further enhance automated scientific survey generation for computer science.
Appendix A Details of the Academic Corpus
@@
Appendix B Social Impact
Automated scientific survey writing brings two-sided social impacts. On the one hand, automated survey generation can save significant time and resources for researchers. This efficiency allows for quick data collection and analysis, enabling more rapid responses to a research domain. Automated survey generation could even provide innovation and lead to discoveries and insights with the help of LLMs. On the other hand, we have shown that the current method for automated survey generation is limited. Such a low-quality survey could poison the academic environment and bring ethical concerns, such as biased content generated by LLMs. Therefore, surveys generated by LLMs should be noted and published in different tracks. Such research direction should be carefully supervised by the collaboration of researchers.
Appendix C Selected Ground Truth Survey
@@
The following three tables provide the complete list of the ground-truth surveys that comprise the SurGE benchmark. In line with our selection methodology outlined in Section 3, these surveys represent high-impact, peer-reviewed works. For each entry, we list its full title, publication year, primary arXiv category, and its citation count. The provided citation counts represent a snapshot taken from Google Scholar on May 10, 2025.

### v2

@@ 7.1 Retrieval-Augmented Generation
Large Language Models (LLMs) are inherently limited by their static, pre-trained parametric knowledge. To address these limitations, Retrieval-Augmented Generation (RAG) has emerged as a key paradigm (Gao et al., 2023; Lewis et al., 2020; Dong et al., 2025; Tu et al., 2025; Su et al., 2025a). By grounding the model in external knowledge, RAG directly addresses several fundamental limitations of LLMs, offering a robust mechanism to mitigate hallucinations (Ji et al., 2023; Su et al., 2024d; b; Wang et al., 2025a), facilitate knowledge updating (Fang et al., 2024a; Wang et al., 2024b; a; 2025b), and enable effective domain adaptation (Yang et al., 2024b; Su et al., 2025c; 2024a; 2024e).
The conventional approach to traditional RAG is built upon the ”Retrieval-then-Read” paradigm (Borgeaud et al., 2022; Guu et al., 2020; Lewis et al., 2020). Within this framework, a user’s query triggers a search module for relevant documents within a large-scale external corpus. This retrieval step is carried out by either an external retriever (Zhai, 2008; Su et al., 2023b; Robertson et al., 2009; Su et al., 2023a; Ma et al., 2023; Fang et al., 2024b) or a more sophisticated retrieval system (Su et al., 2023c; Salemi & Zamani, 2024; Chen et al., 2022). Building upon this foundation, recent work has proposed more advanced RAG architectures to improve efficiency and effectiveness. For instance, Dynamic RAG (Jiang et al., 2022; Su et al., 2024c; Yao et al., 2024) moves beyond a single retrieval step by adaptively triggering the retriever during generation, specifically when the LLM is uncertain during the generation process. From another angle, GraphRAG (Edge et al., 2024) enhances the knowledge source by querying pre-constructed knowledge graphs instead of unstructured text, allowing it to retrieve interconnected facts and relationships. Furthermore, the Parametric RAG paradigm  (Su et al., 2025b; Tan et al., 2025; Fleshman & Van Durme, 2025) alters the knowledge injection step by directly injecting retrieved knowledge into the LLM’s parameters.
@@ 7.2 Long-Form Text Generation and Evaluation
Long-form text generation is substantially more challenging than short-text generation due to its inherent requirements for sustained coherence and rich contextual understanding. Early approaches mainly used generative adversarial networks and reinforcement learning to conduct long-sequence generation (Guo et al., 2018). More recently, large language models have emerged as a strong tool for this task, offering advanced capabilities to handle long-text generation. For example, structured planning techniques and specialized inference mechanisms are proposed (Sloan et al., 2024; Jin et al., 2024) to generate consistent and high-quality clinical reports. Similarly, hierarchical planning frameworks have demonstrated that content control and multi-constraint instruction following can significantly enhance logical flow and topic coverage (Hu et al., 2022; Pham et al., 2024). Beyond medical or other task-specific applications, context-driven retrieval strategies, such as tree-structured retrieval, can support open-domain long-text generation by guiding the model through extensive knowledge sources (Roy et al., 2024). The effective evaluation framework is vital for measuring the quality, factualness, and user-centric utility of the long-form text generation task. Traditional metrics, designed for shorter texts, often fail to capture the intricacies of longer outputs. Recent work has introduced task-focused benchmarks that emphasize user-oriented objectives, such as personalized writing or domain-specific content generation (Kumar et al., 2024; Salemi et al., 2025). In parallel, factuality assessment has attracted growing interest, with methods proposed to evaluate both verifiable and unverifiable claims. Metrics such as VERISCORE and FACTSCORE break down generated text into atomic facts, checking each for consistency against reliable sources (Song et al., 2024; Min et al., 2023). Beyond factual correctness, coherence and structural quality have been studied extensively. Benchmarks like LongGenBench and HelloBench underscore the importance of evaluating a model’s ability to maintain logical organization and clarity over extended passages (Wu et al., 2025; Que et al., 2024).
@@ 7.3 Survey Generation
In the domain of scientific writing, survey generation involves distilling extensive textual resources into a coherent and structured overview. Recent advances in AI-assisted systems have provided prompting-based approaches to expedite the drafting process while preserving content accuracy (Gero et al., 2022; Kacena et al., 2024). One of the most commonly used approaches is retrieval-augmented generation, which combines large-scale knowledge retrieved from documents with language generation empowered by LLMs to yield factually comprehensive overviews (Lewis et al., 2021). Retrieval-augmented generation is often initiated with dense retrieval methods based on dual-encoder architectures to identify highly relevant documents (Karpukhin et al., 2020). Once these documents are retrieved, summarization techniques—spanning top-down, bottom-up, and graph-based ranking methods—play a pivotal role in producing concise yet faithful summaries (Nayeem & Rafiei, 2024; Pang et al., 2022; Bleiweiss, 2023). Building on these retrieval and summarization-based methodologies, automated literature survey generation has garnered increasing attention (Tian et al., 2024a; Lai et al., 2024a). However, existing techniques depend on limited ground truths and employ coarse evaluation metrics, resulting in oversimplified assessments of survey quality (Wang et al., 2024d). To address these challenges, we present a refined ground truth and a multi-dimensional evaluation framework that emphasizes both accuracy and structural coherence. By evaluating quality through multiple dimensions, our proposed framework advances the capabilities of automated survey generation, offering a more comprehensive and rigorous approach to summarizing scientific literature.
@@ 8 Conclusion
In this paper, we introduce SurGE, a comprehensive benchmark designed to address the critical need for standardized and reproducible evaluation in automated scientific survey generation. SurGE provides a large-scale academic corpus, a set of expert-written ground-truth surveys, and a fully automated framework to evaluate surveys on comprehensiveness, citation accuracy, structure, and content. Our experiments reveal significant limitations in state-of-the-art LLM-based systems, highlighting challenges such as incomplete topic coverage and reference hallucination. We believe SurGE will catalyze future research and guide the development of more effective LLM-based systems at the intersection of information retrieval and generative AI for this important task.
9 Ethics Statement
The SurGE benchmark aims to facilitate the advancement of automated survey generation and evaluation in the context of scientific research. By utilizing publicly available metadata from arXiv, our work prioritizes transparency and open access. We do not host or redistribute arXiv’s PDF content, ensuring compliance with arXiv’s Terms of Use and the CC0 public-domain dedication for metadata. While the use of large-scale academic data in SurGE poses minimal ethical risks, we emphasize the importance of ethical data usage, privacy, and responsible development of AI systems. We encourage the research community to consider these aspects when applying or building upon SurGE.
10 Reproducibility Statement
To ensure reproducibility, all code, data, and models for the SurGE benchmark are open-sourced and available at our official GitHub Repository [^5] https://github.com/oneal2000/SurGE. Detailed descriptions of the dataset construction, processing steps, and evaluation methods can be found at our official GitHub Repository and appendices. Our open-access approach is designed to facilitate the replication of results and encourage future research in the field of survey generation using large language models.
Appendix A Reference Extraction and Processing
[…]
Appendix B Details of the Academic Corpus
@@
Appendix C The Use of Large Language Models
Large language models were used as a tool to help refine the language and improve the clarity of this manuscript. Following any such use, the authors reviewed and edited the content to ensure its accuracy and originality, and take full responsibility for the final text.
Appendix D Baseline Implementation
D.1 Retrieval-Augmented Generation
D.2 AutoSurvey
AutoSurvey (Wang et al., 2024d) implements a multi-stage survey generation pipeline that starts with a high-level outline and proceeds through iterative expansions. We adapt it to use our fine-tuned retriever in place of its original retrieval mechanism and keep the number of retrieved references consistent for fairness. In the adapted workflow, we first issue a query based on the topic $t$ to retrieve an initial collection of papers $P_{\text{init}}$. The LLM is then prompted to create a structured outline, which includes main sections and subsections tailored to the subject matter. Next, each section is expanded by conditioning on the subset of papers most relevant to that specific section, producing a draft that includes references in a bracketed format (e.g., “[id]”). Once each section is drafted, the LLM refines it to address factual inconsistencies, stylistic mismatches, and reference-formatting issues. Finally, all refined sections are merged into a coherent final survey, with transitions and citation references carefully aligned. The workflow iterates over these stages, leading to incremental improvements in both thematic coverage and presentation quality.
D.3 StepSurvey
StepSurvey (Lai et al., 2024b) is a more granular generation strategy that also begins with retrieving the top 100 candidate papers for a given topic $t$ but proceeds through distinct planning and drafting phases in a sequential manner. This baseline is proposed by a team named “ID” in the NLPCC2024 competition task 6 (Tian et al., 2024b). Rather than producing an overarching outline at once, it starts by proposing a survey title and a set of primary headings that collectively capture the central themes of the retrieved literature. Subsequently, it uses the primary headings to guide the selection of secondary or finer-grained topics, each mapped to a relevant subset of the retrieved papers. The LLM then produces a full draft by writing each subsection with explicit attention to references and academic conventions, thereby encouraging greater control and consistency across sections. Throughout this process, the system attempts to maintain a balanced level of detail, striving for a clear exposition of important subtopics while avoiding excessive verbosity or redundancy. By structuring the content in incremental steps, StepSurvey aims to achieve coherent organization and thorough coverage of the literature.
Appendix E Prompt Template
This section provides the detailed prompt templates used for our LLM-as-a-Judge evaluation, as described in the main body of the paper. These prompts are specifically designed for GPT-4o to assess the Structure Quality Score (SQS) and the Content Quality of the generated surveys.
Prompt Template for Structure Quality Score You are an AI evaluator. Your task is to compare the generated titles with the target titles and assign a score from 0 to 5 based on their similarity in structure, meaning, and wording. Target Titles: {Ground Truth Titles} Generated Titles: {Generated Titles} Scoring Criteria: 5 – Almost Identical: Nearly all key words match exactly. The meaning is fully preserved. The phrasing and structure are identical or differ only in trivial ways. 4 – Very Similar: Most key words match. The meaning is nearly identical. The phrasing and structure are very close, with minor rewording. 3 – Similar: Several key words are shared. The meaning is largely the same with slight variations. The structure is somewhat similar, but there may be word substitutions. 2 – Somewhat Similar: Some key words are shared, but others are different. The general topic is the same, but the emphasis may differ. The sentence structures are different but not entirely unrelated. 1 – Somewhat Different: Few words overlap, but they are not key terms. The meaning is somewhat related but mostly different. The sentence structures are significantly different. 0 – Completely Different: Nearly no words in common. Completely different meanings. No similarity in structure or phrasing. Instruction: Analyze the generated titles based on the criteria above and provide a single score between 0 and 5.
Prompt Template for Content Quality Score You are an advanced AI language evaluator. Your task is to assess the logical coherence and clarity of the text based on the following criteria: 1. Fluency and Coherence – Does the text flow naturally? Are the sentences well-connected and easy to read? 2. Logical Clarity – Is the reasoning clear and structured? Does the argument progress logically without contradictions? 3. Avoidance of Redundancy – Does the text avoid unnecessary repetition? 4. Clarity of Description – Are ideas, concepts, or events described in a way that is easy to understand? 5. Absence of Errors – Does the text contain grammatical mistakes, spelling errors, or factual inconsistencies? You will provide a score from 0 to 5 based on the following criteria: 5 – Excellent: A score of 5 is awarded to texts that are highly fluent, featuring smooth transitions and a natural flow. The logical progression is clear, well-structured, and easy to follow. There is no redundancy; each sentence contributes meaningfully to the overall message. Furthermore, descriptions are precise and unambiguous, and the text is free of any spelling, grammatical, or factual errors. 4 – Good A score of 4 indicates a text that is mostly fluent but may have minor awkward transitions. Its logical progression is clear, though it might contain slight inconsistencies. There may be some minor redundancy or repetition present. While descriptions are mostly clear, they could contain minor ambiguities, and the text has very few spelling or grammatical errors. 3 – Average A score of 3 applies to texts that are understandable despite containing noticeable awkward phrasing. The logical flow is inconsistent, which may cause some points to feel out of place. Some redundancy or repetition is present and slightly affects readability. Additionally, certain descriptions are vague or unclear, and the text contains some spelling or grammatical mistakes but remains readable. 2 – Poor A score of 2 is assigned when a text is difficult to read due to an awkward structure and poor fluency. Logical inconsistencies make the central argument unclear, and repetitive phrases render the content tedious. Descriptions are vague, making it hard to understand key points, and the submission is characterized by multiple grammatical and spelling errors. 1 – Very Poor A score of 1 denotes a text that is highly disjointed, making it very hard to read. The logical flow is almost nonexistent, with abrupt topic shifts throughout. Redundant sentences are included but add no value to the content. Descriptions are confusing or overly vague, and the text suffers from frequent spelling and grammatical mistakes. 0 – Incoherent A score of 0 is reserved for a text that is completely nonsensical or unreadable. It demonstrates no logical progression or coherence. The content exhibits extreme redundancy or devolves into word salad, and severe errors throughout make it impossible to understand any intended meaning. Instruction Now evaluate the following paragraph based on the criteria above and provide a score from 0 to 5. Paragraph: {Paragraph}
Appendix F Selected Ground Truth Survey
@@
The following three tables provide the complete list of the ground-truth surveys in the SurGE benchmark. For each entry, we list the title, publication year, primary arXiv category, and citation count. The provided citation counts represent a snapshot from Google Scholar on May 10, 2025.